Inc., has sales of $21 million, total assets of $19.1 million, and total debt of $6.8 llon. Assume the profit margin is 8 percent a. What is the companys net income? (Do not round intermediate calculations. Enter b. What is the companys ROA? (Do not round intermediate calculations. Enter your c. What is the companys ROE? (Do not round intermediate calculations. Enter your your answer in dollars not in millions,e.g., 1,234,567) answer as a percent rounded to 2 decimal places, e.g., 32.16.) answer as a percent rounded to 2 decimal places, e.g., 32.16) a. Net income . ROA C. ROE Prev 1of 10 İİİ Next 25 MacBook Air
Answer #1
Net income Sales×Profit margin $ 1,680,000 =21000000*8%
ROA Net income/Total assets 8.80% =1680000/19100000
ROE Net income/Total equity 13.66% =1680000/(19100000-6800000)